OpenStudy (anonymous):

If P divided by 5 =Q, then P divided by 10 = what???

OpenStudy (anonymous):

\[\frac{P}{5}=Q\] \[\frac{P}{10}=\frac{Q}{2}\]

OpenStudy (anonymous):

can write \[\frac{P}{10}=\frac{P}{5}\times \frac{1}{2}=Q\times \frac{1}{2}=\frac{Q}{2}\]